We have a poster board that we must fit squares of size 1 in x 1 in. The poster is 25 in x 18 in. How many squares do you use? This is a 5th grade project. I am 11 years. Thank you!!
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

perimeter = 2(25) + 2(18) = 86 inches

On the other hand, number of squares along border: 2(25 - 2) + 2(18 - 2) + 4 = 82 squares.

Answer by ikleyn(52946)

.

To get the answer, simply multiply the given dimensions


    25 * 18 = 450 square inches


and 450 squares of 1 by 1 inch each.

Solved.